OpenStudy (anonymous):

Two angles are complementary. One of the angles is twice the other. The larger angle has a measure of 120° 60° 45°

OpenStudy (anonymous):

what is it a quadrilateral or triangle ?

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

what do you mean ? The data is not complete ? anyways before you seek help , first tell me wht is a complimentary angle, based on your answer I will guide you forward

OpenStudy (anonymous):

thats all the questions asks. it does not give me more information =/ im assuming triangle?

OpenStudy (anonymous):

good, so lets say its a triangle, then the sum of complimentary angles is 90

OpenStudy (anonymous):

but then the third angle has to be 90

OpenStudy (anonymous):

so now say that one of the angles is x then the other angle is 2x Because it is twice the other so x+2x+90=180 sovle for x

OpenStudy (anonymous):

you shall have your answer

OpenStudy (anonymous):

does that help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now can you tell me the answer? want to see if you get the problem

OpenStudy (anonymous):

what happens if the problem was a quadrilateral

OpenStudy (anonymous):

th answer would be 90 right?

OpenStudy (anonymous):

since a triangle equals 180

OpenStudy (anonymous):

sorry , but the answer is 60 x+2x+90=180

OpenStudy (anonymous):

Oh brain fart!!! I knew that LOL

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

3x=90, solving we get x=30, and the larger angles of the two is 2x = 2*30= 60

OpenStudy (anonymous):

how many sides does a quadrilateral have ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

good, and what is the sum of all angles?

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

great

OpenStudy (anonymous):

now lets say two angles are complimentary , i.e they add up to 90,

OpenStudy (anonymous):

meaning the other two add up to 270

OpenStudy (anonymous):

makes sense. 90+90+90+90

OpenStudy (anonymous):

nope not true,!

OpenStudy (anonymous):

oh. lol yea thats not 360

OpenStudy (anonymous):

x+2x+a+b=360

OpenStudy (anonymous):

and we have no information about relation between a and b

OpenStudy (anonymous):

Dang, thank you so much lol. I'm finally understanding!

OpenStudy (anonymous):

Can you help me with another problem by chance?

OpenStudy (anonymous):